OpenStudy (imtiaz7):

The expression ax^3 - 9x^2 + bx +3a is exactly divisible by x^2 -5x +6. Find the value of a and b. @mary.rojas

OpenStudy (anonymous):

write ax^3-9x^2+bx+3a=(x^2-5x+6)(ax+p) since the coeff of the x^3 term is a. p is a new unknown. compare the coeffs of x^2,x and constants. You'll get 3 equations and 3 unknowns

OpenStudy (imtiaz7):

try to solve @kelvinltr

OpenStudy (anonymous):

a=-2 from constants 3a=6p which means a=2p for x^2 terms -9=p-5a try the rest